B= F/(P-V)
in which
B= units sold to break even point
F=fixed costs
P= price per unit
V= variable costs
Network systems is introducing a new network card. Suppose Network systems knos its fixed costs are $600,000, its vaariable costs are $500 per card, and it must sell 15,000 cards to break even the first year. What is the minimum price per unit it should charge?

Answer by checkley75(3666)

15000=F/(P-V)
15000=600000/(500-V) CROSSS MULTIPLY
15000(500-V)=600000
7500000-15000V=600000
-15000V=600000-7500000
-15000V=-6900000
V=-6900000/-15000
V=460 PRICE PER CARD TO BREAK EVEN.